OpenStudy (anonymous):

solve: sqrt{x+3}-3=x

OpenStudy (kinggeorge):

You have \[\sqrt{x+3}-3=x\implies\sqrt{x+3}=x+3\]

OpenStudy (kinggeorge):

Which numbers \(y\) exist such that \[\sqrt{y}=y?\]

OpenStudy (anonymous):

so x=-3??

OpenStudy (kinggeorge):

That's one solution. There is one other.

OpenStudy (anonymous):

-2 ?

OpenStudy (kinggeorge):

Perfect.